# CarDreamer: Open-Source Learning Platform for World Model based Autonomous Driving

### v1

@@ 5.1 World Model Training & Evaluation
We use a small DreamerV3 model with only 18M parameters Figure 4 as the model backbone. A small DreamerV3 has 32 CNN multiplier, 512 GRU and MLP units, and the MLP has only two layers within its RSSM [10]. The small memory overhead is around 10GB which allows us to train on a single NVIDIA 4090 GPU alongside running CARLA simulator.
We train the agent on each task. The change in reward curves with respect to time steps is shown in Figure 2. Simpler tasks with less traffic, such as ’right turn simple’ and ’lane merge’, typically converge within 50k steps (about 1 hour), whereas tasks involving denser, aggressive traffic flows, which require collision avoidance, take approximately 150k-200k steps to converge (about 3 to 4 hours).
@@
  • Average Distance (m): Represents the average distance traveled by the ego vehicle across all episodes before the episode terminates, either through task completion or due to a failure such as a collision or timeout.
@@
  • Waypoint Distance: This metric quantifies the average divergence from the desired route waypoints. It assesses the vehicle’s ability to adhere to the planned path, reflecting its navigation accuracy and precision in following the given trajectory.
@@
Figure 3: Sampled images during one episode when completing different tasks.
@@
Tasks Success Rate Avg. Distance (m) Collision Rate Avg. Speed (m/s) Wpt. Distance
[…]
@@ 5.2 Predictions in Different Observation Modalities
The results, illustrated in Fig.4, compare the ground-truth images with the imagined ones across three modalities. The first row displays the ground truth observation images, the second row the WM’s imagined outcomes, and the third row the differences between them. We selected frames within an imagination horizon of up to 64 time steps.
@@ 5.3 Benefits of V2V Communication
A distinctive feature of CarDreamer is its ability to facilitate easy customization of the level to which vehicles communicate. Vehicles can share FOV views, leading to different observability. Besides, they can even share intentions (represented by vehicles’ planned waypoints) for better planning. We utilize this feature to evaluate the impact of communication. An agent is trained and tested on the “right turn hard” task under different settings, i.e., different obsrevability and whether it has access to others’ intentions. The “right turn hard” task is particularly suitable for testing observability and intention communication due to the dense traffic and frequent potential for collisions from vehicles outside the FOV.
The reward curves are shown in Figure 5 and some performance metrics are shown in Table 2. Note that the successful behavior in making the right turn is approximately indicated by rewards exceeding 250 in our reward functions. The results show that limited observability or lack of intention sharing impedes the agent from completing the task. The evenly sampled images during one episode (shown in Figure 6) provides a good explanation: the agent adopts a conservative and sub-optimal policy where it stops at the crossroad to avoid collision. For example, in the first three rows of Figure 6, the agent stops moving before merging into the car flow. In contrast, the complete information enables the ego vehicle to successfully execute the right turn.
(a) Observability.
(b) Intention sharing.
Figure 5: Reward curves in different communication settings.
Table 2: Metrics in different communication settings.
Settings Success Rate Avg. Distance (m) Collision Rate Avg. Speed (m/s) Wpt. Distance
[…]
Figure 6: Sampled images during one episode in different communication settings.
@@ 6 Conclusion
We introduced CarDreamer, an open-source learning platform tailored for the development and evaluation of WM based RL algorithms in autonomous driving. CarDreamer offers a comprehensive set of built-in tasks, a flexible task development suite, and an integrated world model backbone, all aimed at facilitating rapid prototyping of driving tasks and algorithm testing within this specialized domain. With its modular design and diverse task configurations, CarDreamer establishes itself as a flexible and challenging testbed for assessing the performance of WM based autonomous driving systems. The experiments we conduct using our platform gives a comprehensive evaluation of the performance of DreamerV3 different driving tasks. We emphasize its predictive accuracy across different observation modalities and the significant impact of communication on performance.

### v2

@@ 5.1 World Model Training & Evaluation
We use a small DreamerV3 model with only 18M parameters as the model backbone. A small DreamerV3 has 32 CNN multipliers, 512 GRU and MLP units, and the MLP has only two layers within its RSSM [10]. The small memory overhead is around 22GB which allows us to train on a single NVIDIA 4090 GPU alongside running CARLA simulator.
We train the agent on each task. The change in reward curves with respect to time steps is shown in Figure 2. All the tasks here are trained with intention sharing and full observability. Simpler tasks with less traffic, such as ‘right turn simple’ and ‘lane merge’, typically converge within 50k steps (about 1 hour), whereas tasks involving denser, aggressive traffic flows, which require collision avoidance, take approximately 150k-200k steps to converge (about 3 to 4 hours).
@@
Figure 3: Sampled images during one episode in different tasks.
@@
Tasks Success Rate Collision Rate Avg. Speed (m/s)
Right turn simple 99.21% $\pm$ 0.49% 0.00 % $\pm$ 0.00% 3.19 $\pm$ 0.01
Right turn medium 98.11 % $\pm$ 0.22% 1.89 % $\pm$ 0.22% 2.99 $\pm$ 0.02
Right turn hard 99.58 % $\pm$ 0.42% 0.42 % $\pm$ 0.42% 2.92 $\pm$ 0.02
Left turn simple 100.00% $\pm$ 0.00% 0.00% $\pm$ 0.00% 3.21 $\pm$ 0.01
Left turn medium 97.58% $\pm$ 0.61% 2.42% $\pm$ 0.61% 3.04 $\pm$ 0.03
Left turn hard 92.36% $\pm$ 3.03% 7.64 % $\pm$ 3.03% 2.97 $\pm$ 0.01
Overtake 100.00% $\pm$ 0.00% 0.00% $\pm$ 0.00% 3.12 $\pm$ 0.03
Four lane 96.83% $\pm$ 3.17% 3.17% $\pm$ 3.17% 3.47 $\pm$ 0.00
Navigation 80.95% $\pm$ 6.41% 7.58% $\pm$ 4.01% 3.97 $\pm$ 0.08
Lane merge 90.61% $\pm$ 1.17% 9.39% $\pm$ 1.17% 5.17 $\pm$ 0.01
Roundabout 96.00% $\pm$ 2.31% 4.00% $\pm$ 2.31% 3.58 $\pm$ 0.00
Traffic lights 92.65% $\pm$ 0.85% 0.00% $\pm$ 0.00% 2.26 $\pm$ 0.03
Stop sign 97.94% $\pm$ 1.09% 0.00% $\pm$ 0.00% 1.80 $\pm$ 0.01
@@ 5.2 Predictions in Different Observation Modalities
The results, illustrated in Figure 4, compare the ground-truth images with the imagined ones across three modalities. The first row displays the ground truth observation images, the second row the WM’s imagined outcomes, and the third row the differences between them. We selected frames within an imagination horizon of up to 64 time steps.
@@ 5.3 Benefits of V2V Communication
A distinctive feature of CarDreamer is its ability to facilitate easy customization of the level to which vehicles communicate. Vehicles can share FOV views, leading to different observability. Besides, they can even share intentions (represented by vehicles’ planned waypoints) for better planning. We utilize this feature to evaluate the impact of communication. An agent is trained and tested on the “right turn hard” task under different settings, i.e., different observability and whether it has access to others’ intentions. The “right turn hard” task is particularly suitable for testing observability and intention communication due to the dense traffic and frequent potential for collisions from vehicles outside the FOV.
The reward curves are shown in Figure 5 and the performance metrics are shown in Table 2 and Table 3. Note that the successful right turn is approximately indicated by the rewards exceeding 300. The results show that limited observability or lack of intention sharing impedes the agent from completing the task. The success rates of tasks with the above limits are significant lower than the task with full information. The evenly sampled images during one episode (shown in Figure 6) provides a good explanation: the agent adopts a conservative and sub-optimal policy. For example, in the ’FOV’ and ’SFOV’ setting in Figure 6, the agent hesitates to move on since they have insufficient information about coming vehicles. In the ’No intention sharing’ setting, the agent cuts in at an inappropriate timing that leads to a collision without shared intentions. In contrast, the complete information enables the ego vehicle to successfully execute the right turn.
Figure 5: Reward curves in different observability settings.
Table 2: Metrics in different observability settings (with intention sharing).
Settings Success Rate Collision Rate Avg. Speed (m/s)
Full Observability 99.58 % $\pm$ 0.42% 0.42 % $\pm$ 0.42% 2.92 $\pm$ 0.02
FOV Observability 12.44 % $\pm$ 1.79% 87.56 % $\pm$ 1.79% 2.78 $\pm$ 0.01
SFOV Observability 29.52 % $\pm$ 3.46% 70.48 % $\pm$ 3.46% 2.86 $\pm$ 0.09
Table 3: Metrics in different intention sharing settings (with full observability).
Settings Success Rate Collision Rate Avg. Speed (m/s)
Intention Sharing 99.58 % $\pm$ 0.42% 0.42 % $\pm$ 0.42% 2.92 $\pm$ 0.02
No Intention Sharing 90.71 % $\pm$ 0.61% 8.94 % $\pm$ 0.58% 3.08 $\pm$ 0.00
Figure 6: Sampled images during one episode in different communication and observability settings.
@@ 6 Conclusion
In this work, we introduced CarDreamer, an open-source learning platform tailored for the development and evaluation of WM based RL algorithms in autonomous driving. CarDreamer offers a comprehensive set of built-in tasks, a flexible task development suite, and an integrated world model backbone, all aimed at facilitating rapid prototyping of driving tasks and algorithm testing within this specialized domain. With its modular design and diverse task configurations, CarDreamer establishes itself as a flexible and challenging testbed for assessing the performance of WM based autonomous driving systems. The experiments we conduct using our platform gives a comprehensive evaluation of the performance of DreamerV3 different driving tasks. We emphasize its predictive accuracy across different observation modalities and the significant impact of communication on performance.